Translate to a system of equations and solve.
The sum of two numbers is 14. Six times the first number minus three times the second number is 3. Find the numbers.

Answer by eperette(173)

x + y = 14
6x - 3y = 3
there are two ways to solve for this, i will use the elimination method:
3( x + y = 14)
6x - 3y = 3
*****************
3x + 3y = 42
6x - 3y = 3
__________________
9x + 0 = 45
9x/9 = 45/9
x = 5
then:
x + y = 14
5 + y = 14
5 - 5 + y = 14 - 5
0 + y = 9
so the numbers are: 5 and 9
check:
5 + 9 = 14
6(5) - 3(9) = 30 - 27 = 3
answer is correct